OpenStudy (anonymous):

What is the value of the coefficient of "a" when the quadratic equation y = (5x − 2)(2x + 3) is written in standard form? −11 −6 7 10

OpenStudy (anonymous):

y = (5x − 2)(2x + 3) \[\huge y = (5x − 2)(2x + 3)\] \[\huge \rightarrow y = 10x^2 + 15x -4x -6\] \[\huge \rightarrow y = 10x^2 + 11x -6\] comparing it with \[\huge y = ax^2 + bx +c\] we find a= 10 Thus last option is correct answer.

OpenStudy (anonymous):

Thanks!!!